OpenStudy (anonymous):

what it means to take square root of a negative number ? and what is the meaning of square root of a positive number ? please give some practical examples for both, like getting length from area of a square piece...

OpenStudy (anonymous):

imaginary number\[\sqrt{-1}=i\]

OpenStudy (anonymous):

Getting length from Area means?? Can you suggest any example to this..??

OpenStudy (anonymous):

if you know the area of a square you can square root that value in order to get the length of a side

OpenStudy (anonymous):

@waterineyes i have edited the question now. pls read now :P

OpenStudy (anonymous):

We cannot take the square root of negative number.. See, there is a little bit difference.. If I say take the square root of 16 Then: \[\sqrt{16} = \pm 4\]

OpenStudy (anonymous):

square root of negative number is an imaginary number

OpenStudy (anonymous):

See if I have an area of Square = 36, Then we know: \[Area = Side^2\] Now I want to side so I will do this as: \[Area = Side^2 \implies 36 = Side^2\] \[Side^2 = 36 \implies Side = \sqrt{36} \implies Side = \pm 6\] Now we get two values of sides... But, a length cannot be negative like -5cm is not possible.. So, \(Side = 6\) will be the answer..

OpenStudy (anonymous):